Jenny is twice as old as Jimmy. Jesse is 10 years older then jimmy. If the sum of their ages is 90, how old is Jenny?
Found 2 solutions by lwsshak3, DrBeeee:
Answer by lwsshak3(11628)

Jenny is twice as old as Jimmy. Jesse is 10 years older then jimmy. If the sum of their ages is 90, how old is Jenny?
***
let x=Jimmy's present age
2x=Jenny's present age
x+10=Jesse's present age
x+2x+x+10=90
4x=80
x=20
2x=40
how old is Jenny? 40

Answer by DrBeeee(684)

To avoid all the J's,
Let a = Jenny's age
Let b = Jimmy's age
Let c = Jesse's age
REM: when you have 3 unknowns, you need 3 equations!
The problem statement provides us with the following three equations.
(1) a = 2*b,
(2) c = 10 + b, and
(3) a + b + c = 90
Put (1) and (2) into (3) to get
(4) 2*b + b + 10 +b = 90 or
(5) 4*b = 90 - 10 or
(6) 4*b = 80 or
(7) b = 20
From (1) we get
(8) a = 2*20 or
(9) a = 40
And from (2) we get
(10) c 10 + 20 or
(11) c = 30
Let's check our ages with (3).
Is (40 + 20 + 30 = 90)?
Is (90 = 90)? Yes
Answer: Jenny is 40 years of age.